I can't figure out how to do question 5 (part 1 and 2) and question 6, could you provide explanationQuestion #4 One way to limit the choices for your unknown is to consider the sign of AHgoln. Considering your results from TaQuestion #6 Given the accepted AH soln of the salts you chose in question 4 listed in the table repeated below, give the lowe
Answer #1

Q5. part 1

For trial 1

The experimental value for \Delta solH can be calculated by the following formula;

Q = m.c. \Delta T; where Q is the heat released or absorbed, m is the mass of water, c is specific heat capacity of water, and  \DeltaT is the change in temperature (Ti - Tf),

\therefore Q = 50.0 g \times 4179.6 J/kg. K \times (21.5 - 27.3) K \times (1kg/1000g) = -1212.084 J

[Note: (i) c for water = 4179.6 J/kg.K; (ii)The kg in 'c' is changed to grams by multiplying with the conversion factor]

Now, \Delta solH in J/g for trial 1 = Q/(mass of solid) = (-1212.084)/(5.029g) = -241.02 J/g

For trial 2

Following exactly the same procedure for trial 2

\DeltasolH in J/g for trial 2 = -254.60 J/g

-----------------------------------------------------------------------------------------------------------

Q5. part 2

\DeltasolH (average) = [\DeltasolH(trial 1) + \Delta solH (trial 2)]/2 = -247.81 J/g

-----------------------------------------------------------------------------------------------------------

Q6.

The salt is clearly Na2CO3 whose given value of \Delta solH = -267 J/g (Theoretical value)

% error = [(Theoretical value - Experimental value or average)/Theoretical value] \times 100 = 7.19 %